OpenStudy (faolan):

Will fan and medal!!! just have one little question.. what is sqrt6 * sqrt2 what is sqrt6 * sqrt3

OpenStudy (mayankdevnani):

\[\large \bf \sqrt{6} \times \sqrt{2}=\sqrt{6*2}\]

OpenStudy (faolan):

sqrt 12?

OpenStudy (mayankdevnani):

yep

OpenStudy (faolan):

ok thx. also theres 6 * sqrt 2

OpenStudy (mayankdevnani):

try your own! i'll check

OpenStudy (faolan):

sqrt 6?

OpenStudy (faolan):

@mayankdevnani

OpenStudy (mayankdevnani):

\[\large \bf \sqrt{6} \times \sqrt{3}\]

OpenStudy (faolan):

thats sqrt 18

OpenStudy (mayankdevnani):

correct

OpenStudy (faolan):

but what would 6 * sqrt2 be?

OpenStudy (mayankdevnani):

\[\large \bf 6 \times \sqrt{2}=6\sqrt{2}\]

OpenStudy (faolan):

really?

OpenStudy (mayankdevnani):

yep

OpenStudy (mayankdevnani):

or i have another way

OpenStudy (mayankdevnani):

\[\large \bf \sqrt{36}=6\] so, \[\large \bf 6\sqrt{2}=\sqrt{36} \times {\sqrt{2}}\]

OpenStudy (faolan):

first one worked for me.......

OpenStudy (faolan):

so would the answer to this be: -6 sqrt2 + 6 sqrt3 Rationalize the denominator and simplify. https://mvusdk12.owschools.com/media/g_alg02_ccss_2014/7/q11168.gif

OpenStudy (mayankdevnani):

go and learn over there

OpenStudy (faolan):

ive been rejected... :'(

OpenStudy (mayankdevnani):

what

OpenStudy (faolan):

ths though I had totally forgotten about that site.....

OpenStudy (mayankdevnani):

so go and learn good luck